Suppose y = mx + b is a math model for actual time as a function of estimated time, where y represents actual time and x represents estimated time and m and b are constants. If m = 2.8 and b = -2.6, find y when x is 30 min.
My answer = 22.72 min
Thanks again!

Answer by Alan3354(69443)

Suppose y = mx + b is a math model for actual time as a function of estimated time, where y represents actual time and x represents estimated time and m and b are constants. If m = 2.8 and b = -2.6, find y when x is 30 min.
My answer = 22.72 min
--------------
y = mx + b
y = 2.8x - 2.6
-----
@ x = 30, y = 2.8*30 - 2.6
y = 81.4